# ACES: Automatic Cohort Extraction System for Event-Stream Datasets

### v2

@@ 2.3 Command-Line Interface Arguments
The Hydra framework (Yadan, 2019) enhances the CLI by enabling flexible run configurations and argument parsing for cohort extractions. For instance, specific arguments are required to define the external source dataset for data loading. Depending on the chosen format (the data.standard argument), either the path to the data file (for meds or direct) or the path to the dataset directory (for esgpt) must be specified to indicate the external source data from which ACES will extract the cohort. Additionally, cohort_dir and cohort_name are essential for locating and loading the task configuration file, as well as for saving results and logging operational data.
@@
An overview of the computational profile of ACES is available in Section 4.1. Additionally, for users dealing with large datasets, ACES can also be run over a collection of sharded files, extracting and storing the matching cohort for each shard individually in matching file paths. This can greatly increase computational efficiency by facilitating the processing of different shards in parallel via Hydra’s multi-run launchers [^3] https://hydra.cc/docs/1.0/plugins/joblib_launcher/.
@@
8 "data.shard=\\$(expand_shards <folder>/<num>)" # Sweeps over shards
@@ 2.4 Extraction Output
Finally, with a dataset configured for predicates and a task configuration, ACES will execute the extraction for the cohort and return a table where each row is a valid instance as per the criteria defined in the configuration file. Hence, each instance can be included in our cohort used for the downstream ML task. At the most basic level, the table contains the patient identifiers of our cohort, a user-defined timestamp that indexes prediction time, and a task label derived from a user-specified predicate. In addition, for each of the interrelated windows, a start and end timestamp is provided to segment the patient record, along with a summary of the number of predicates evaluated in said window. Please see Figure 3 for a sample output table.
2.5 Python Usage
In addition to the command-line tool, we also provide a Python interface to allow researchers to easily leverage ACES for cohort extraction in their code pipelines. A full tutorial is provided at: https://eventstreamaces.readthedocs.io/en/latest/notebooks/tutorial_meds.html. Please also refer to the PDF compilation of the tutorial notebook in the Supplementary Material (Pages 31-36).
3 ACES Technical Design
Key Design Principles
ACES is built with the following key design principles in mind:
  1. 1.
    Structured Medical Data as Event-Stream Data: ACES views structured medical data as “event-stream” data1, enabling the extraction of meaningful tasks and cohorts from continuous patient records.
  2. 2.
    Compatibility with Any Dataset: ACES is designed to be compatible with any dataset, provided the necessary dataset-specific predicates are extracted and formatted as an event-stream (this step is automated for any dataset that can be expressed in the MEDS (Arnrich et al., 2024) or ESGPT (McDermott et al., 2023) formats) [^4] Note: This includes datasets that can be converted into these formats using existing ETLs, such as OMOP datasets through the MEDS OMOP ETL..
  3. 3.
    Separation of Configuration Components: ACES clearly separates dataset-specific configuration components from dataset-agnostic configuration components, ensuring flexibility and ease-of-use.
  4. 4.
    Simple Configuration Language: ACES leverages a simple configuration language to maximize portability and facilitate community feedback and iteration on ML tasks and environments.
  5. 5.
    Parallelization for Computational Efficiency: ACES relies on parallelization over patient shards to handle large datasets efficiently, ensuring scalability and performance.
Algorithm Overivew
The key novel principle behind the ACES algorithm is that we decompose the task of determining whether a section of a patient’s record meets certain criteria specified in a configuration file into a series of recursive analyses. These analyses evaluate whether specific windows, defined relative to the realizations of other windows in the patient record, meet their criteria. By imposing only minimal constraints on the structure of the task configuration file, we can ensure that tasks can always be extracted using this recursive approach, thus forming the foundation of the entire ACES system. For more details, please see the Appendix and the technical documentation in the Supplementary Materials.
Figure 3: Sample output table from ACES. Continuing the example from Figure 2, we obtain this table after executing ACES extraction. The subject_id, index_timestamp, and label form the foundation of the defined cohort and are what the user would need, at a minimum, to perform downstream tasks. In addition to these three columns, there exists a column for each of the defined windows in the task configuration, with its value being a structured object containing information about that particular window, such as the start and end times, as well as a summary of predicate counts. As seen in the values of this table, the timestamp of the end of the input window is always 24 hours after the trigger timestamp for each sample (row), and contains more than 5 of _ANY_EVENT—exactly matching our previously defined task configuration.
4 Using ACES: A Repository of Task Configuration Examples
@@
To demonstrate the flexibility and utility of ACES, we define and publicly release the task configurations described in Table 1, both with dataset-agnostic criteria and with dataset-specific predicate realizations based on the MEDS version of the public MIMIC-IV dataset. Additional usage profiles, tasks, and areas, as well as details of the task cohorts extracted from MIMIC-IV, can be found in the Appendix and documentation. These various tasks have been previously studied, and ACES will facilitate their conceptual reproducibility to encourage benchmarking efforts and ensure robustness in ML for healthcare.
@@
Table 1: A collection of sample configuration files for various common predictive tasks on MIMIC-IV (Johnson et al., 2023). These tasks can be easily generalized to other datasets, such as e-ICU (Pollard et al., 2018) or other private intensive care unit (ICU) and inpatient datasets by simply swapping out appropriate predicate definitions. pastecode.io links have been provided.
@@ 4.1 Computational Profile
To establish an overview of the computational profile of ACES, a collection of various common tasks was queried on MIMIC-IV (Johnson et al., 2023) in the MEDS format. The MIMIC-IV MEDS schema has approximately 50,000 patients per shard with an average of approximately 80,500,000 total event rows per shard over seven shards.
Table 2: Performance statistics for various common predictive tasks on a single MEDS shard of MIMIC-IV (Johnson et al., 2023), including the number of patients, the number of samples, total runtime (in seconds), and maximum memory usage (in MiBs).
@@
All experiments were executed on a Linux server with 36 cores and 340 GBs of RAM available. A single MEDS shard was used to provide a bounded computational overview of ACES, which is applicable even when scaled to larger datasets using Hydra3. For instance, if one shard costs $M$ memory and $T$ time, then $N$ shards may be executed in parallel with $N*M$ memory and $T$ time, or in series with $M$ memory and $T*N$ time.
@@ 5.1 Additional Related Work
In addition to the existing tools discussed in Section 1, there are several other areas of related work relevant to ACES. Firstly, ACES could be directly connected with and support various health CDMs, such as OMOP, FHIR, PCORNET, and i2b2 (OHDSI, 2021; Bender & Sartipi, 2013; PCORnet, 2023; Murphy et al., 2010), rather than through ETLs. These models provide already-accepted standardized frameworks for organizing and analyzing healthcare data, and integrating ACES directly with them could greatly enhance its utility and interoperability. Similarly, frameworks such as DescEmb (Hur et al., 2022) and GenHPF (Hur et al., 2024) hold great synergistic potential with ACES, and we believe that they can be complementary in enabling new kinds of cross-dataset training, transfer learning, and evaluation. Static benchmarks that provide standardized datasets, evaluation metrics, and baseline methods for a range of clinical problems, such as YAIB (van de Water et al., 2024), Harutyunyan et al.’s multitask learning clinical prediction benchmarks (Harutyunyan et al., 2019), and EHR-PT (McDermott et al., 2021a), can also be directly integrated with ACES to facilitate robust ML in healthcare. Lastly, ACES can be used in conjunction with various health data management tools, such as ESGPT (McDermott et al., 2023), TemporAI (Saveliev & van der Schaar, 2023), PyHealth (Yang et al., 2023), and OMOPLearn (Kodialam et al., 2021). These tools offer functionalities for pre-processing, managing, and analyzing health data for downstream tasks, and integrating ACES with them directly can streamline ML workflows.
Beyond healthcare, ACES is applicable to data from a variety of other domains, such as for financial, climate, or social media data—essentially, ACES could be used for any structured, longitudinal data that can be reformatted as an event-stream. This versatility makes ACES a powerful library for extracting and analyzing complex event-based datasets across different fields.
@@ 5.2 Limitations & Future Roadmap
ACES has several key limitations that can be addressed in future work. Firstly, while already very expressive, ACES’s task configuration language can still be further expanded. Expressing more complex kinds of predicates, window aggregations, labeling functions, and criteria would expand the scope of ACES significantly. ACES also seeks to provide direct support for cohort extraction based on unstructured data (notes and memos) in the future, though currently, such predicates need to be manually extracted by the user.
@@
ACES is also very well poised to capture more complex patterns of task and cohort relationships, including prescribed systems of case/control matching, automated bias analyses, or propensity re-weighting over excluded populations. It is also possible to enable users to nest ACES configuration files to leverage extracted task labels as new predicates in more complex ACES extraction processes.
@@
To enhance the scalability of ACES, we will ensure that the expanding MEDS standard will always be supported. Direct interoperability with existing resources in this space, in particular ATLAS and OHDSI vocabulary-derived cohort definitions, is a high priority area for future work.
@@
Finally, with the standardization that ACES offers, new opportunities for human interaction with data are also made available, such as via a natural language interface to define ACES predicates or configuration files and, thus, to extract downstream tasks, patient cohorts, or derived datasets in a code-free manner on diverse input EHR formats. We aim to also explore the viability of leveraging large language models (LLMs) to directly format predicate and criteria definitions given a data dictionary; preliminary work on this front shows very promising results!
5.3 ACES for Enabling a New Kind of Benchmark
@@
In addition to the clear impact of ACES on reproducibility, robustness, and accessibility of health datasets and ML for healthcare, we also feel that ACES is critical for a “new kind of benchmark” in the field—and, in so being, is a portent of what needs to come should ML for healthcare progress to a more productive, communal, and impactful stage.
@@
In particular, we argue that for this field to progress in the manner desired by the community, and most likely to be maximally positively impactful for all patients, we need to develop methodologies to test, share, and develop ML solutions across diverse datasets in a meaningful and reproducible manner, even without said datasets being publicly available to general researchers. This capability is critical because, without it, we will never be able to offer new inductive insights about which methods are most likely to work best on novel, private data. In other words, if we cannot test our model training recipes across the diverse sets of clinical care settings, populations, and conceptual dataset schemas that exist in the real world, we similarly cannot expect those training recipes to generalize to said set of myriad downstream deployment areas.
@@
Libraries like ACES, which make it as easy as possible for users to share the conceptual definitions of their tasks and prediction areas across datasets—in such a way that their colleagues can use them even over independent, private datasets—can help transform the kinds of benchmarking studies we can perform in ML for healthcare towards those that permit generalizable assessment of ML training recipes across datasets, clinical areas, and more.
@@ 6 Conclusion
In this work, we present the Automatic Cohort Extraction System for Event-Stream Datasets (ACES). ACES is a system designed to intuitively define cohorts and downstream tasks of interest for learning representations and reliably and automatically extract those cohorts from arbitrary datasets that are in an event-stream format. This system enables significantly greater shareability of task definitions, reproducibility of ML training and evaluation recipes, and is as easy to use as installing a package via pip and running a simple command-line tool. We feel that ACES will be integral in the development of new kinds of benchmarks in ML for healthcare, which can be explored across both public and private datasets alike, and characterize the populations and tasks of interest in a manner that cleanly separates dataset-specific components from shareable dataset-agnostic components. To learn more about ACES and use it today in your work, please visit our GitHub repository at: https://github.com/justin13601/aces, and the online ACES Documentation at: https://eventstreamaces.readthedocs.io/en/latest/index.html.
7 Video Demonstration
To complement this paper and the documentation, we have provided a very short online video demonstration of ACES in action, available here: https://youtu.be/i_hCaHDydqA. It overviews the Python API and walks through a running example of extracting a cohort for an in-hospital mortality prediction task from a synthetic MEDS dataset. Additionally, it briefly covers several simple and complex real-world task configuration files.
8 Reproducibility Statement
We have provided links to the ACES library source code and pastecode.io snippets of configuration files used in the experiments. Reproducible tutorial notebooks and technical implementation details are included in the Supplementary Materials (Pages 31-36 and 41-51, respectively) and are accessible via the online public documentation at: https://eventstreamaces.readthedocs.io/en/latest/.
9 Acknowledgments and Disclosure of Funding
MBAM gratefully acknowledges support from a Berkowitz Postdoctoral Fellowship at Harvard Medical School. JX greatly appreciates support from supervisors David W. Eyre (Big Data Institute, University of Oxford) and Curtis P. Langlotz (Center for Artificial Intelligence in Medicine & Imaging, Stanford University). We also acknowledge contributions from Tom Pollard (Massachusetts Institute of Technology).
Appendix A Documentation
The full ACES Online Documentation is available at: https://eventstreamaces.readthedocs.io/en/latest/. We have also included a compiled PDF version of this documentation in the Supplementary Material.
To answer specific questions about ACES, please see the below index (links to the external online documentation, as well as references to chapters and page numbers of the ancillary PDF documentation are provided):
@@ Appendix A Documentation
    Sample task descriptions and specifications are provided on the Task Examples page of the online documentation (https://eventstreamaces.readthedocs.io/en/latest/notebooks/examples.html). Please refer to Chapter 3, Pages 21-26.
@@
      For an overview of predicates and how they form the foundation of ACES, please refer to the Predicates DataFrame section (https://eventstreamaces.readthedocs.io/en/latest/notebooks/predicates.html) in Chapter 4, Pages 27-30.
@@
      A window in ACES is defined as a segment in the patient record (https://eventstreamaces.readthedocs.io/en/latest/algorithm.html#window). For details on how to define a window (https://eventstreamaces.readthedocs.io/en/latest/readme.html#windows), please refer to Chapter 1.3.3, Page 9.
@@
    For general ACES usage instructions (https://eventstreamaces.readthedocs.io/en/latest/usage.html#quick-start), please refer to Chapter 2.1, Pages 13-16. Additionally, brief end-to-end instructions (https://eventstreamaces.readthedocs.io/en/latest/readme.html#instructions-for-use) are also available in Chapters 1.2 and 1.3, Pages 4-10.
@@
      For detailed instructions on using ACES CLI (https://eventstreamaces.readthedocs.io/en/latest/usage.html#detailed-instructions), please refer to the Usage Guide in Chapter 2.2, Pages 16-20.
@@
      For a step-by-step tutorial on using the ACES Python API (https://eventstreamaces.readthedocs.io/en/latest/notebooks/tutorial_meds.html), please refer to the Code Example Notebook in Chapter 5, Pages 31-41.
@@
    ACES with & vs. Other Tools
    For an overview of how ACES could be used with other existing complementary tools for reproducible machine learning (https://eventstreamaces.readthedocs.io/en/latest/readme.html#complementary-tools), please refer to Chapter 1.4.2, Page 10.
    For an overview of ACES and other existing alternative tools for semi- or fully-automated cohort extraction (https://eventstreamaces.readthedocs.io/en/latest/readme.html#alternative-tools), please refer to Chapter 1.4.3, Page 11.
@@
    For technical details on the ACES configuration language (https://eventstreamaces.readthedocs.io/en/latest/configuration.html), please refer to the Configuration Language Specification section in Chapter 6.1, Pages 41-44.
@@
    For a glossary of terminology used throughout ACES (https://eventstreamaces.readthedocs.io/en/latest/technical.html#algorithm-terminology), please refer to the Algorithm Terminology section in Chapter 6.3, Pages 46-47.
@@
    For technical details on the ACES algorithm (https://eventstreamaces.readthedocs.io/en/latest/technical.html#algorithm-design), please refer to the Algorithm Design section in Chapter 6.4, Pages 48-51.
@@
    For the complete ACES module documentation (https://eventstreamaces.readthedocs.io/en/latest/api/modules.html), including doctests that ensure algorithm correctness, please refer to the Module API sections in Chapter 8, Pages 55-125.
@@
    For an overview of the computational profile of ACES (https://eventstreamaces.readthedocs.io/en/latest/profiling.html), please refer to the Computational Profile section in Chapter 7, Page 53.
@@
    For additional examples of configuration files and criteria of different machine learning for health tasks, please refer to the MEDS-DEV benchmarking effort on GitHub (https://github.com/mmcdermott/MEDS-DEV/tree/main).

### v3

@@ 2.4 Command-Line Interface
The Hydra framework (Yadan, 2019) enhances the CLI by enabling flexible run configurations and argument parsing for cohort extractions. For instance, specific arguments are required to define the external source dataset for data loading. Depending on the chosen format (the data.standard argument), either the path to the data file (for meds or direct) or the path to the dataset directory (for esgpt) must be specified to indicate the external source data from which ACES will extract the cohort. Additionally, cohort_dir and cohort_name are essential for locating and loading the task configuration file, as well as for results and operational logging.
@@
An overview of the computational profile of ACES is available in Section 3.1. Additionally, for users dealing with large datasets, ACES can also be run over a collection of sharded files, extracting and storing the matching cohort for each shard individually in corresponding file paths. This can greatly increase computational efficiency by facilitating the processing of different shards in parallel via Hydra’s multi-run launchers [^2] https://hydra.cc/docs/1.0/plugins/joblib_launcher/.
Sample command for cohort extraction over multiple MEDS shards using ACES CLI and Hydra:
@@
8 data.shard="\$(expand_shards <folder>/<num>)" # Sweeps over shards
Figure 4: Example configuration file for the binary prediction of in-hospital mortality 48 hours after admission. References to predicates and windows are italicized and bolded, respectively. (A) Dataset-specific task predicates. These concepts are needed to conceptually capture this task and are used as constraints and boundaries for windows of the patient record. For instance, in this example, the value of “$ADMISSION$” denotes a hospital admission event in the source dataset. (B) A window of the task specifying the task inputs for downstream models. Suppose we’d like to use all historic patient data up to and including 24 hours past the admission. We could also place an arbitrary criterion requiring more than 5 records in this window to ensure that the extracted cohort contains sufficient input data. (C) Trigger events for the task, which are hospital admissions as we’d like to make a mortality prediction for each admission. (D) A window of the task specifying a gap in the patient timeline. Suppose we’d like to set a minimum length of admission for our cohort (e.g., 48 hours). A temporal constraint (minimum window duration) of 48 hours could then be set to represent this requirement. (E) A window of the task specifying the task target, which is set from the end of (D) to the immediately subsequent $discharge$ or $death$ predicate. This creates our binary label classes for the task (i.e., $discharge=0$; $death=1$). All windows are interrelated on the patient timeline, as shown by how each window references another in the configuration file.
2.5 Python API
In addition to the command-line tool, we also provide a Python interface to allow researchers to easily leverage ACES for cohort extraction in their deep learning code pipelines. A full tutorial is available on the ACES online documentation.
@@ 2.6 Extraction Output
Finally, with a dataset configured for predicates and a task configuration file, ACES will execute the extraction for the cohort and return a table where each row is a valid instance as per the criteria defined in the configuration file. Hence, each instance can be included in our cohort used for the downstream ML task. At the most basic level, the table contains the patient identifiers of our cohort, a user-defined timestamp that indexes prediction time, and a task label derived from a user-specified predicate. In addition, for each of the interrelated windows, a start and end timestamp is provided to segment the patient record, along with a summary of the number of predicates evaluated in that window.
3 Using ACES: A Repository of Example Task Configuration Files
@@
To demonstrate the flexibility and utility of ACES, we define and publicly release the task configuration files described in Table 1, both with dataset-agnostic criteria and with dataset-specific predicate realizations based on the MEDS version of the public MIMIC-IV dataset. These various tasks have been previously studied, and ACES will facilitate their conceptual reproducibility to encourage benchmarking efforts and ensure robustness in ML for healthcare.
@@
Table 1: A collection of sample configuration files for various common predictive tasks on MIMIC-IV. These tasks can be easily generalized to other datasets, such as e-ICU or other private intensive care unit (ICU) and inpatient datasets by simply swapping out appropriate predicate definitions.
@@ 3.1 Computational Profile
To establish an overview of the computational profile of ACES, the collection of tasks from Table 1 was extracted on MIMIC-IV. The MIMIC-IV MEDS schema has approximately 50,000 patients per shard with an average of approximately 80,500,000 total event rows per shard over seven shards. However, only a single shard was used to provide the bounded computational overview of ACES in Table 2, as the results are applicable even when scaled to larger datasets using Hydra2. For instance, if one shard costs $M$ memory and $T$ time, then $N$ shards may be executed in parallel with about $N*M$ memory and $T$ time, or in series with about $M$ memory and $T*N$ time.
Table 2: Performance statistics for various common predictive tasks on a single MEDS shard of MIMIC-IV. This particular shard was 5,494.39 MiBs on disk as a parquet file and 9,661.41 MiBs when loaded in memory as a dataframe. It consisted of 80,301,208 rows corresponding to 47,954 unique patients. All experiments were executed on a Linux server with 36 cores and 340 GBs of RAM available.
@@ 4.1 Additional Related Work
In addition to the existing tools discussed in Section 1, there are several other areas of related work relevant to ACES. Firstly, ACES serves as a middle ground between solutions that focus on specific CDMs, such as OHDSI’s ATLAS Gold et al. (2024) and i2b2’s PIC-SURE (Stedman et al., 2024). Compared to these tools, ACES balances capability with greater ease of use and improved communicative value. ACES is also not tied to a particular CDM. Built on a flexible event-stream format, ACES is a no-code solution with a descriptive input format, permitting easy and wide iteration over task definitions. It can be applied to a variety of schemas, making it a versatile tool suitable for diverse research needs. ACES could also be directly connected with existing health CDMs through ETLs, such as OMOP (Reich et al., 2024), i2b2 (Murphy et al., 2010), FHIR (Bender & Sartipi, 2013), and PCORnet (Fleurence et al., 2014). These models provide already-accepted standardized frameworks for organizing and analyzing healthcare data, and supporting them could greatly enhance the utility and interoperability of ACES. Similarly, frameworks such as DescEmb (Hur et al., 2022) and GenHPF (Hur et al., 2024) hold great synergistic potential with ACES, and we believe that they can be complementary in enabling new kinds of cross-dataset training, transfer learning, and evaluation. Static benchmarks that provide standardized datasets, metrics, and baseline methods for a range of clinical problems, such as YAIB (van de Water et al., 2024), multitask learning clinical prediction benchmarks (Harutyunyan et al., 2019), and EHR-PT (McDermott et al., 2021a), can also be directly integrated with ACES to facilitate robust ML in healthcare. Lastly, ACES can be used in conjunction with various health data management tools, such as TemporAI (Saveliev & van der Schaar, 2023), PyHealth (Yang et al., 2023), OMOP-learn (Kodialam et al., 2021), and DPM360 (Suryanarayanan et al., 2021). These tools offer functionalities for pre-processing, managing, and analyzing health data for downstream tasks, and integrating ACES with them directly can streamline ML workflows.
Beyond healthcare, ACES is applicable to data from a variety of other domains, such as for finance, climate, or social media data — essentially, ACES could be used for any structured, longitudinal data that can be reformatted as an event-stream. This versatility makes ACES a powerful library for extracting and analyzing complex event-based datasets across different fields.
@@ 4.2 Limitations & Future Roadmap
ACES has limitations that can be addressed in future work. Firstly, while already very expressive, the ACES task configuration language can still be further expanded. Expressing more complex kinds of predicates, window aggregations, labeling functions, and criteria would expand the scope of ACES significantly. ACES also seeks to provide direct support for cohort extraction based on unstructured data (notes and memos) in the future. Currently, such predicates need to be manually extracted by the user, but with the help of community contributions, we hope to be able to incorporate automatic feature extraction from clinical notes, or even images, and integrate them into configuration files for cohort extraction.
@@
ACES is also very well poised to capture more complex patterns of task and cohort relationships, including prescribed systems of case-control matching, automated bias analyses, or propensity re-weighting over excluded populations. It is also possible to enable users to nest ACES configuration files to leverage extracted task labels as new predicates in more complex tasks and querying processes.
@@
To enhance the scalability of ACES, we will seek to maintain support for the expanding MEDS standard. Direct interoperability with other existing resources in this space, in particular ATLAS and its OHDSI vocabulary-derived cohort definitions, is a high priority area for future work.
@@
Finally, with the standardization that ACES offers, new opportunities for human interaction with data are also made available, such as via a natural language interface to define ACES predicates or configuration files and, thus, to extract downstream tasks, patient cohorts, or derived datasets in a code-free manner on diverse input EHR formats. We aim to explore the viability of leveraging large language models (LLMs) to directly format predicate and criteria definitions given a data dictionary, and to automatically construct configuration files from natural language.
4.3 ACES as a Catalyst for a New Era of Benchmarking
@@
In addition to the clear impact of ACES on reproducibility, robustness, and accessibility of ML for healthcare, we also feel that ACES is critical for a “new kind of benchmark” in the field — and, in so being, is a portent of what needs to come should ML for healthcare progress to a more productive, communal, and impactful stage (McDermott, 2025).
@@
In particular, we argue that for this field to progress in the manner desired by the community and to maximize positive impact for all patients, we need to develop methodologies to test, share, and develop ML solutions across diverse datasets in a meaningful and reproducible manner, even without said datasets being publicly available to general researchers. This capability is critical because, without it, we will never be able to offer new inductive insights about which methods are most likely to work best on novel, private data. In other words, if we cannot test our model training recipes across the diverse sets of clinical care settings, populations, and conceptual dataset schemas that exist in the real world, we similarly cannot expect those training recipes to generalize well to a myriad of downstream deployment areas.
@@
Libraries like ACES, which make it as easy as possible for users to share the conceptual definitions of their tasks and prediction areas across datasets — in such a way that their colleagues can use them even over independent, private datasets — can help transform the kinds of benchmarking studies that we can perform in ML for healthcare. For instance, the MEDS Decentralized Extensible Validation (MEDS-DEV) (McDermott et al., 2024) effort is one such step towards enabling the generalizable assessment of ML training recipes across datasets, clinical areas, and beyond.
@@ 5 Conclusion
In this work, we present the Automatic Cohort Extraction System for Event-Stream Datasets (ACES). ACES is a system designed to intuitively define cohorts and downstream tasks of interest for representation learning and reliably extract those cohorts from arbitrary datasets in event-stream formats. This system enables significantly greater shareability of task definitions, reproducibility of ML training and evaluation recipes, and is as easy to use as installing a package via pip and running a simple command-line tool. We feel that ACES will be integral in the development of new kinds of benchmarks in ML for healthcare, which can be explored across both public and private datasets alike, as well as help characterize populations and tasks of interest in a manner that cleanly separates dataset-specific components from shareable dataset-agnostic components. To learn more about ACES and use it today in your work, please visit our GitHub repository at: https://github.com/justin13601/aces, and the ACES online documentation at: https://eventstreamaces.readthedocs.io/en/latest.
Acknowledgments and Disclosure of Funding
MBAM gratefully acknowledges support from a Berkowitz Postdoctoral Fellowship at Harvard Medical School. JG is funded by the National Institutes of Health through NIH-USA R01CA294033. JX greatly appreciates support from supervisors David Eyre (University of Oxford) and Curtis Langlotz (Stanford University). We also acknowledge valuable contributions by Tom Pollard (Massachusetts Institute of Technology) and by the broader MEDS ecosystem of contributors and users.
Appendix A ACES Online Documentation Index
The full ACES online documentation is available at: https://eventstreamaces.readthedocs.io/en/latest. We have also included a compiled PDF version of this documentation in the Supplementary Material.
To answer specific questions about ACES, please see the below index for the PDF documentation (links to the associated online documentation are also provided).
@@ Appendix A ACES Online Documentation Index
    Sample task descriptions and specifications are provided in the Task Examples section in Chapter 3 (https://eventstreamaces.readthedocs.io/en/latest/notebooks/examples.html).
@@
      For an overview of predicates and how they form the foundation of ACES, please refer to the Predicates DataFrame section in Chapter 4 (https://eventstreamaces.readthedocs.io/en/latest/notebooks/predicates.html).
@@
      A window in ACES represents a segment in the patient record. For details on how to define a window, please refer to Chapter 1.3.3 (https://eventstreamaces.readthedocs.io/en/latest/readme.html#windows).
@@
    For general ACES usage instructions, please refer to Chapter 2.1 (https://eventstreamaces.readthedocs.io/en/latest/usage.html#quick-start). Additionally, brief end-to-end instructions are also available in Chapters 1.2 and 1.3.
@@
      For detailed instructions on using ACES CLI, please refer to the Usage Guide in Chapter 2.2 (https://eventstreamaces.readthedocs.io/en/latest/usage.html#detailed-instructions).
@@
      For a step-by-step tutorial on using the ACES Python API, please refer to the Code Example Notebook in Chapter 5 (https://eventstreamaces.readthedocs.io/en/latest/notebooks/tutorial_meds.html).
@@
    ACES with / vs. Other Tools
    For an overview of how ACES could be used with other existing complementary tools for reproducible ML, please refer to Chapter 1.4.2 (https://eventstreamaces.readthedocs.io/en/latest/readme.html#complementary-tools).
    For an overview of how ACES compares to other existing alternative tools for semi- or fully-automated cohort extraction, please refer to Chapter 1.4.3 (https://eventstreamaces.readthedocs.io/en/latest/readme.html#alternative-tools).
@@
    For technical details on the ACES configuration language, please refer to the Configuration Language Specification section in Chapter 6.1 (https://eventstreamaces.readthedocs.io/en/latest/technical.html#configuration-language-specification).
@@
    For a glossary of terminology used throughout ACES, please refer to the Algorithm Terminology section in Chapter 6.3 (https://eventstreamaces.readthedocs.io/en/latest/technical.html#algorithm-terminology).
@@
    For technical details on the ACES algorithm, please refer to the Algorithm Design section in Chapter 6.4 (https://eventstreamaces.readthedocs.io/en/latest/technical.html#algorithm-design).
@@
    For the complete ACES module documentation, including doctests that ensure algorithm correctness, please refer to the Module API sections in Chapter 8 (https://eventstreamaces.readthedocs.io/en/latest/api/modules.html).
@@
    For an overview of the computational profile of ACES, please refer to the Computational Profile section in Chapter 7 (https://eventstreamaces.readthedocs.io/en/latest/profiling.html).
@@
    For additional examples of configuration files and criteria of different ML for healthcare tasks, please refer to the MEDS-DEV benchmarking effort on GitHub: https://github.com/mmcdermott/MEDS-DEV.
Appendix B Comparative Experiments
We conducted preliminary experiments to quantitatively compare ACES with other alternative tools. Using the OMOP version of the MIMIC-IV Demo, as well as a synthetic dataset of 1,000 patients generated using Synthea [Walonoski et al., 2017] and converted into OMOP, we queried four tasks using ACES and two of the most comparable tools, OMOP-learn and DPM360.
We collected metrics including script runtime, peak memory usage (in MiBs), lines of code required (including configuration files and any template code needed to execute extraction), and human time spent. All experiments were conducted on a default A100 GPU instance with 84 GB of RAM and 12 CPU cores from Google Cloud Platform’s Compute Engine.
Table B.1: Quantitative comparison of ACES and other comparable cohort extraction tools across datasets and tasks.
Dataset Method Task Runtime (s) Peak Memory (MiB) Lines of Code Human Time (s)
Synthea-1000 ACES via MEDS First 24h in-hospital mortality 0.386 389 35 120
30d post-hospital-discharge mortality 0.236 351 32 90
30d re-admission 0.337 355 22 60
End-of-life prediction 0.449 421 28 120
DPM360 via OMOP First 24h in-hospital mortality 5.932 390 205 2,126
30d post-hospital-discharge mortality 4.188 550 257 1,200
30d re-admission 6.260 870 288 2,020
End-of-life prediction 4.901 387 222 1,500
MIMIC-IV Demo ACES via MEDS First 24h in-hospital mortality 0.617 545 35 180
30d post-hospital-discharge mortality 0.301 509 32 90
30d re-admission 0.455 532 22 90
End-of-life prediction 0.349 589 28 300
OMOP-learn via OMOP First 24h in-hospital mortality 12.220 688 172 3,623
30d post-hospital-discharge mortality 8.608 587 199 2,441
30d re-admission 19.710 640 168 2,998
End-of-life prediction 24.540 932 251 12,000
We also qualitatively compared the adaptability of these approaches to any other given ML task. While ACES requires simple modifications to the task configuration file to capture new task logic or cohort criteria, new ATLAS executions, bespoke SQL queries, and changes to Python parameters may be needed for DPM360 and OMOP-learn.
While we acknowledge potential biases in these results due to our familiarity with ACES and only surface exposure to other extraction tools, we have aimed for a fair and objective evaluation. Based on our experience and preliminary user feedback, we find ACES to be intuitive and believe it offers a significantly improved workflow with more comprehensive functionality.